OpenStudy (anonymous):

Lydia bought a shirt at 20% off its retail price of $40. She paid 5% tax on the price after the discount. How much did Lydia pay for the shirt? $42 $36.60 $34 $33.60

OpenStudy (jebonna):

So first thing to do is to find 20% of $40. You can do this by doing this: 40/100 x 20 = $8 This means 20% of $40 = $8 Now take 20% away from the $40: 40-8 = $32 after discount Now find 5% of $32 doing this: 32/100 x 5 = $1.6 So add on the 5% tax to $32 = 32 + 1.6 = $33.6 So in the end the shirt costs $33.6 I hope this helps :)

OpenStudy (anonymous):

thanks

OpenStudy (jebonna):

Its ok :)